OpenStudy (anonymous):

Add x√3x + √27x^3. (all of 27x^3 is in the sqrt sign)

OpenStudy (solomonzelman):

\(\Large\color{blue}{ \bf x\sqrt{3x} +\sqrt{27x^{3} } }\) \(\Large\color{blue}{ \bf x\sqrt{3x} +x\sqrt{27x } }\) \(\Large\color{blue}{ \bf x\sqrt{3x} +x\sqrt{9\times 3 x } }\) \(\Large\color{blue}{ \bf x\sqrt{3x} +3x\sqrt{3x } }\) \(\Large\color{blue}{ \bf 4x\sqrt{3x } }\)

OpenStudy (anonymous):

Thanks so much!